OpenStudy (anonymous):

what is 2y+4= -x in mx+b form?

OpenStudy (anonymous):

y=mx+b 2y=-x-4

OpenStudy (anonymous):

Then divide by 2

OpenStudy (anonymous):

y=-1x/2-2

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

@mjmiller1 I think you meant to put y=-1/2x-2 not y=-1x/2-2.

OpenStudy (texaschic101):

@RoseDryer is correct....y = - 1/2x - 2 :)